Sonja Cc is a small firm that makes outfits for customers tailored for their functions, in May 2021 she was ordered for a baby shower (BS)

The details for the Job BS are as follows:

Direct materials: 12 units at N$124.

Direct labour: 20 hours for department A at N$25 per hour.

10 hours for department B at N$30 per hour

12 hours for department C at N$27.50 per hour


The firm allocates manufacturing overheads budgeted at N$2500 in department A, N$2000 in department B and N$3000 in department C.

The firm expects a profit of 55% on the total cost of BS


Required:

a)Calculate the cost of the baby shower BS

b)Calculate the selling price of a baby shower (BS)

c)If the number of outfits are the same for each type of function, which function should Sonja choose to produce and why?


1
Expert's answer
2022-03-25T15:07:21-0400

a)Calculate the cost of the baby shower BS

"124\times12 =1488"

"20\times25 =500"

"10\times30 = 300"

"12\times27.50 =330"

The total variable cost "= N2618"


The fixed cost is ;

"2500+2000+3000 =7500"


Total Cost "= 7500+2618"

"= N10118"


b) The selling price of the BS


"55" percent profit is

"\frac{155}{100}\times 10118 = 15682.9"


Per unit = "\frac{15682.9}{12}= N1306.9"